How do i solve the following problem?
2-i)^3

Answer by rothauserc(4718)

(2-i)^3 = (2-i)(2-i)(2-i)
= (4-4i+i^2)(2-i)
note that i^2 = -1
= (3-4i)(2-i)
= 6-11i+4i^2
= 2-11i
note that i = square root(-1)
= 2 - 11*square root(-1)